# Crop Row Switching for Vision-Based Navigation: A Comprehensive Approach for Efficient Crop Field Navigation

### v1

@@ Abstract
Vision-based mobile robot navigation systems in arable fields are mostly limited to in-row navigation. The process of switching from one crop row to the next in such systems is often aided by GNSS sensors or multiple camera setups. This paper presents a novel vision-based crop row-switching algorithm that enables a mobile robot to navigate an entire field of arable crops using a single front-mounted camera. The proposed row-switching manoeuvre uses deep learning-based RGB image segmentation and depth data to detect the end of the crop row, and re-entry point to the next crop row which would be used in a multi-state row switching pipeline. Each state of this pipeline use visual feedback or wheel odometry of the robot to successfully navigate towards the next crop row. The proposed crop row navigation pipeline was tested in a real sugar beet field containing crop rows with discontinuities, varying light levels, shadows and irregular headland surfaces. The robot could successfully exit from one crop row and re-enter the next crop row using the proposed pipeline with absolute median errors averaging at 19.25 cm and 6.77°for linear and rotational steps of the proposed manoeuvre.
I Introduction
@@
The integration of robotics and autonomous systems in agriculture has enabled precision farming operations leading to effective use of time and resources. Achieving autonomous navigation in agricultural robots is an enabling technology that must be optimised for the deployment of autonomous robots for precision agriculture. The existing autonomous navigation solutions for in-field navigation, albeit being efficient, often relies on expensive sensors such as Real-Time Kinematic Global Navigation Satellite System (RTK-GNSS) sensors. Camera-based agricultural robot navigation systems are a popular alternative to these expensive sensors [1]. Implementation of such vision systems is often limited to crop row following behaviour [2]. In most such systems, row switching is achieved by the aid of GNSS sensors or multiple cameras to identify row end and re-entry to the next row [1, 3, 4].
Figure 1: Crop row switching state machine. $d_{r}$: Distance between current and next crop row, A: Initial Detection of the EOR, B: Robot is at the EOR, C: Robot traversed a distance equal to its length into the headland, D: Robot turn 90°towards the next row direction from state C, E: Robot traverse $d_{r}$ distance forward from state D, F: Robot turn 90°towards the next row direction from state E, G: Robot re-enter the next crop row.
@@
The premise of vision-based navigation in agricultural robots is to reduce the cost of the overall robotic system with the aim of increased adoption of such technologies. To this end, a vision-based navigation system that uses a single front-mounted camera to perform in-row navigation and row switching would be in line with this objective. In our previous work, we have developed a vision-based crop row detection algorithm [5] and an in-row navigation framework [6] that then uses the crop row detection to guide the robot through a single crop row only based on RGB images. The crop row switching algorithm presented in this paper relates to our previous work on vision-based crop row detection [5] and navigation [6] in arable fields. The newly proposed algorithm serves as a vital bridge, seamlessly connecting with the established in-row navigation algorithm to dexterously integrate into a comprehensive, fully autonomous field-scale navigation behaviour. This row-switching algorithm could also integrated seamlessly with any other existing in-row navigation methods to eliminate the need for GNSS sensors in row-switching. This existing system can follow a crop row based on RGB image input and it can also identify the location of the end of crop rows when the robot is reaching towards the headland area. The crop row switching algorithm presented is be triggered upon the detection of the end of row (EOR) and navigates the robot towards the entry point of the next crop row to be traversed.
@@
  • A vision-based re-entry point detection algorithm that identifies the relative 3D position of the entry point of the adjacent crop row.
@@
  • A comprehensive evaluation and validation of an autonomous navigation system in a real arable field with a mobile robot deployed the proposed system.
II Related Work
@@
Research into vision-based navigation systems for crop row following is an extensively explored subject area [7, 8]. Existing robot navigation technologies for agri-robotics research include GNSS, inertial navigation systems (INS) and light detection and ranging (LiDAR) [9]. Each of these existing systems presents a cost-benefit trade-off which results in a lack of reliable navigation options for an affordable solution in crop row navigation. The vision-based infield navigation technologies available for arable field navigation explores the usage of RGB and depth images to identify the crop rows [1, 10]. The existing systems for such infield navigation mainly uses image segmentation, object detection or image matching methods to identify the crop rows for robot navigation [7]. Despite the considerable interest in vision-based crop row following systems, the presence of reliable vision-based crop row switching algorithms remains limited [2]. Most of the existing vision-based navigation systems depend completely or partially on GNSS, INS or Lidar-based solutions for crop row switching rather than a fully vision-based solution [7]. The crop row switching algorithms proposed in [2, 11, 12] completely rely on the RTK-GNSS-based sensors to perform the crop row switching manoeuvre. However, GNSS-based systems are not considered a simple and straightforward solution for agricultural robot navigation since they also need multiple redundancies in place for effective operation due to multi-path reflections and signal blockage[13]. Some systems also use manual control to perform the headland turn [14]. Fully vision-based solutions in agricultural robot navigation depend on multiple cameras on the robot for maintaining the localisation during the row switching process [15, 16]. The methods that use a single camera also demand special requirements such as variable field of view [17].
@@
End of row (EOR) detection is an important step for any crop row switching algorithm since it serves as the starting point for any crop row switching manoeuvre [6]. The EOR detection scheme proposed in [18] employs image binarisation using classic computer vision methods to calculate the pixel count in binary masks to determine the EOR. However, the pixel count thresholds may vary depending on the stage of the plant growth and these thresholds must be matched to each growth stage when used for the long term. The pixel counts of binary image representing the crop would not yield the spatial localisation of EOR within a given image, while it only generates an EOR trigger based on the image. In contrast, the EOR detection method proposed by authors of [19] uses $Cr$ channel of $YCbCr$ colour space to calculate the position of the EOR within a given image. Such methods pose the advantage of early detection and potential failures due to noisy images in the method presented in [18]. However, such colour-based EOR detection methods are highly susceptible to distortions caused by external field variations. The advantages of deep learning-based methods in EOR detection outperform such colour-based methods [5, 6]. A 3D point cloud-based row-end detection was introduced in [14] where the EOR is identified by detection of height drop between the plant and the ground within the point cloud. This approach is mostly limited to crops with noticeable height differences relative to the ground level or crops at later growth stages. The novel EOR detection algorithm presented in this paper is based on the deep learning-based crop row detection model from our earlier work [5], considering the limitations of the existing EOR detection methods. Identification of the relative distance between two crop rows is also vital for accurate crop row switching. The distance between two crop rows often referred to as ”inter-row distance” is considered as a fixed distance in existing crop row switching methods [3, 20]. The relative position between the current robot pose and the next crop row could vary due to imperfections in planting or due to a slight offset during robot navigation. Therefore, active perception of the relative position of the re-entry point to the next crop row to be traversed is a useful attribute. The re-entry point detection algorithm proposed in this work estimates the relative distance to the next crop row based on crop row segmentation mask from Triangle Scan Method (TSM) [5] and depth data. Such active perception of inter-row space helps to eliminate any potential row-switching failures caused by varying inter-row space or inaccurate positioning of the robotic platform within the currently traversing crop row. Four most common headland turning patterns are semi-circular turn, U-turn, light bulb turn (a.k.a $\Omega$ turn) and switch-back turn [11]. The semi circular turn describes a half circle with a constant radius while U-turn describes two quarter circle turns (90°) centered with a linear traversal stage. These methods are typically used in smaller robots with tighter turning radii relative to the inter-row distance [18, 15, 20]. The $\Omega$ turn and the switch-back turn are mostly used in robots with constrained maneuverability such as Ackerman steering robots or tractors [11, 21, 22]. Considering the skid-steering robot configuration of the robotic platform used in this work, U-turn pattern was chosen to execute the row switching manoeuvre.
III Methodology

### v2

# A Vision-Based Navigation System for Arable Fields

@@ Abstract
Vision-based navigation systems in arable fields are an underexplored area in agricultural robot navigation. Vision systems deployed in arable fields face challenges such as fluctuating weed density, varying illumination levels, growth stages and crop row irregularities. Current solutions are often crop-specific and aimed to address limited individual conditions such as illumination or weed density. Moreover, the scarcity of comprehensive datasets hinders the development of generalised machine learning systems for navigating these fields. This paper proposes a suite of deep learning-based perception algorithms using affordable vision sensors for vision-based navigation in arable fields. Initially, a comprehensive dataset that captures the intricacies of multiple crop seasons, various crop types, and a range of field variations was compiled. Next, this study delves into the creation of robust infield perception algorithms capable of accurately detecting crop rows under diverse conditions such as different growth stages, weed density, and varying illumination. Further, it investigates the integration of crop row following with vision-based crop row switching for efficient field-scale navigation. The proposed infield navigation system was tested in commercial arable fields traversing a total distance of 4.5 km with average heading and cross-track errors of 1.24 and 3.32 cm respectively.
keywords:
vision-based navigation, autonomous systems, agricultural robots, robotic vision, row following, arable fields
1 Introduction
The global agricultural sector is at a critical juncture amidst the increasing food demand due rising global population, labour shortages, environmental sustainability and the drive for maximising the efficiency of food production with precision agricultural practices. Addressing the economic impacts of labour shortages and inefficient agricultural practices is simultaneously important while achieving sustainable food production minimising adverse environmental impact. Therefore, exploring technological solutions that mitigate these labour and economic challenges while remaining acutely conscious of their environmental impact, ensuring a harmonious balance between productivity and ecological preservation should be inherent in future agricultural robotic solutions [5]. The state-of-the-art (SOTA) solutions in agricultural robotic navigation void of these characteristics due to higher costs and poor reliability accommodates the need to develop cheaper vision-based solutions to fill in the demands of future agricultural automation. The integration of robotics and autonomous systems in agriculture has enabled precision farming operations leading to effective use of time and resources.
@@ 1 Introduction
Achieving autonomous navigation in agricultural robots is an enabling technology that must be optimised for the deployment of autonomous robots for precision agriculture. The existing autonomous navigation solutions for in-field navigation, albeit being efficient, often rely on expensive sensors such as Real-Time Kinematic Global Navigation Satellite System (RTK-GNSS) sensors. Camera-based agricultural robot navigation systems are a popular alternative to these expensive sensors [6]. Implementation of such vision systems is often limited to crop row following behaviour [11]. In most such systems, row switching is achieved by the aid of GNSS sensors or multiple cameras to identify row end and re-entry to the next row [6, 2, 13]. Identification of the initial turning direction and the last crop row to be traversed while navigating an entire field is also important for this class of navigation algorithms to become self-reliant. Both of these problems entail the detection of the robot being at the crop row next to the edge of the field. The initial turn direction and the last row detection are often not discussed in the existing field scale navigation systems. Most existing methods require the user to define these parameters by declaring the initial turn direction and number of crop rows.
@@
The premise of vision-based navigation in agricultural robots is to reduce the cost of the overall robotic system with the aim of increased adoption of such technologies. To this end, a vision-based navigation system that uses a single front-mounted camera to perform in-row navigation and row switching would be in line with this objective. In our previous work, we have developed a vision-based crop row detection algorithm [28] and an in-row navigation framework [27] that then uses the crop row detection to guide the robot through a single crop row only based on RGB images. The crop row switching algorithm presented in this paper relates to our previous work on vision-based crop row detection [28] and navigation [27] in arable fields. The newly proposed crop row switching algorithm serves as a vital bridge, seamlessly connecting with the established in-row navigation algorithm to dexterously integrate into a comprehensive, fully autonomous field-scale navigation behaviour. This row-switching algorithm could also integrated seamlessly with any other existing in-row navigation methods to eliminate the need for GNSS sensors in row-switching. This existing system can follow a crop row based on RGB image input and it can also identify the location of the end of crop rows when the robot is reaching towards the headland area. The crop row switching algorithm presented is be triggered upon the detection of the end of row (EOR) and navigates the robot towards the entry point of the next crop row to be traversed.
A complete field scale navigation system could be realised by facilitating the row-following behaviour with a complementary crop row switching algorithm to enable the headland traversal to switch between adjacent crop rows during infield navigation. Existing vision-based methods of crop row switching in arable fields require multiple cameras [33] and symmetric robotic setups [1]. Certain existing approaches demand hybrid vision and GNSS solutions where row switching behaviour is dependent on the GNSS-based navigation [32]. The lack of a vision-only single camera-based solution hinders the cost-effective nature of vision-based navigation systems with the addition of multiple specialised camera setups and GNSS sensors. The initial row switching direction must be manually specified in the existing methods. The existence of autonomous field coverage direction identification is also an important feature of field scale navigation algorithms which is attributed to a user configuration parameter in the existing systems [6]. Such improvements would be highly valued at the adoption of such agricultural robotic systems by the end users where the system functions fully autonomously without needing to configure to each specific deployment instance. We have developed a field-scale crop row navigation system by integrating our vision based in-row navigation methods [28, 27] with the crop row switching and infield orientation algorithms introduced in this paper.
@@
  • A fully vision-based infield navigation system that can perform field scale navigation in crop row fields.
@@
  • An vision-based infield orientation algorithm that identifies the initial turning direction for the robot when reaching the end of first crop row in a field.
@@
  • A comprehensive evaluation and validation of the autonomous infield navigation system in simulation and a real arable field with a mobile robot deployed with the proposed system.
The remainder of this paper is arranged as follows: Section 2 discusses the existing approaches and their shortcomings. Section 3 introduces the vision-based navigation paradigm in arable fields explaining the steps followed during infield navigation of the robot. Section 4 summarises the three experiments conducted to verify the efficacy of the proposed system. Section 5 concludes the outcomes of the proposed work suggesting future developments needed to optimize the proposed system.
2 Related Work
@@
Research into vision-based navigation systems for crop row following is an extensively explored subject area [31, 26]. Existing robot navigation technologies for agri-robotics research include GNSS, inertial navigation systems (INS) and light detection and ranging (LiDAR) [19]. Each of these existing systems presents a cost-benefit trade-off which results in a lack of reliable navigation options for an affordable solution in crop row navigation. The vision-based infield navigation technologies available for arable field navigation explores the usage of RGB and depth images to identify the crop rows [6, 3]. The existing systems for such infield navigation mainly uses image segmentation, object detection or image matching methods to identify the crop rows for robot navigation [31]. Majority of the existing studies on vision-based infield navigation focuses on the row following aspect of infield navigation [20, 7]. Relatively less work has been done towards enabling vision-based crop row switching and headland turning for infield navigation [31]. Switching from one crop row to another often remains unsolved using computer vision [32]. The vision-based crop row switching system demands several perception capabilities including end of row (EOR) detection, re-entry point identification and localisation within the headland [13, 31, 16].
@@
End of row (EOR) detection is an important step for any crop row switching algorithm since it serves as the starting point for any crop row switching manoeuvre [27]. EOR detection was implemented in the vision-based cotton row following algorithm by Li et al. [16]. However, their system requires a driver to take over the control during the row-switching stage. This method of manual row-switching behaviour within autonomous row-following systems is a repeating trend in several other infield navigation schemes [36]. A vision-based EOR detection algorithm was developed based on the percentage of vegetation pixels in the image [4]. This method is limited to the headland without any vegetation while it will fail to detect verdant headlands based on a complete vegetation detection-based approach. They also use GNSS-based navigation to execute the row switching behaviour which would require the robot to include a GNSS receiver. The low-cost GNSS receiver used in [14] robot caused the system to guide the robot towards an undesired crop row due to limited GNSS accuracy. An accurate RTK-GPS system would be required for such GNSS-based systems to perform crop row switching with high fidelity. Some systems rely on GNSS to locate the end of row position along with other sensors such as IMU and compass [13, 35]. The EOR detection scheme proposed in [12] employs image binarisation using classic computer vision methods to calculate the pixel count in binary masks to determine the EOR. However, the pixel count thresholds may vary depending on the stage of the plant growth and these thresholds must be matched to each growth stage when used for the long term. The pixel counts of a binary image representing the crop would not yield the spatial localisation of EOR within a given image, while it only generates an EOR trigger based on the image. In contrast, the EOR detection method proposed by authors of [36] uses $Cr$ channel of $YCbCr$ colour space to calculate the position of the EOR within a given image. Such methods pose the advantage of early detection and potential failures due to noisy images in the method presented in [12]. However, such colour-based EOR detection methods are highly susceptible to distortions caused by external field variations. The advantages of deep learning-based methods in EOR detection outperform such colour-based methods [28, 27]. Lidar and ultrasonic ranging sensors are also used for EOR detection, particularly in the vineyard and orchard navigation scenarios where the absence of adjacent tree rows indicate EOR [25, 29]. A 3D point cloud-based row-end detection was introduced in [14] where the EOR is identified by detection of height drop between the plant and the ground within the point cloud. This approach is mostly limited to crops with noticeable height differences relative to the ground level or crops at later growth stages. The EOR detection algorithm referenced throughout this paper is introduced from our previous work [27], considering the limitations of the existing EOR detection methods.
Identification of the relative distance between two crop rows is also vital for accurate crop row switching. The distance between two crop rows often referred to as "inter-row distance" is considered as a fixed distance in existing crop row switching methods [2, 9]. The relative position between the current robot pose and the next crop row could vary due to imperfections in planting or due to a slight offset during robot navigation. Therefore, active perception of the relative position of the re-entry point to the next crop row to be traversed is a useful attribute. The re-entry point detection algorithm proposed in this work estimates the relative distance to the next crop row based on crop row segmentation mask from Triangle Scan Method (TSM) [28] and depth data. Such active perception of inter-row space helps to eliminate any potential row-switching failures caused by varying inter-row space or inaccurate positioning of the robotic platform within the currently traversing crop row.
Despite the considerable interest in vision-based crop row following systems, the presence of reliable vision-based crop row switching algorithms remains limited [11]. Most of the existing vision-based navigation systems depend completely or partially on GNSS, INS or Lidar-based solutions for crop row switching rather than a fully vision-based solution [31]. The crop row switching algorithms proposed in [11, 10, 21] completely rely on the RTK-GNSS-based sensors to perform the crop row switching manoeuvre. However, GNSS-based systems are not considered a simple and straightforward solution for agricultural robot navigation since they also need multiple redundancies in place for effective operation due to multi-path reflections and signal blockage[24]. Some systems also use manual control to perform the headland turn [14]. Fully vision-based solutions in agricultural robot navigation depend on multiple cameras on the robot for maintaining the localisation during the row switching process [1, 17]. The methods that use a single camera also demand special requirements such as variable field of view [34]. Infield navigation algorithms that completely rely on vision sensors use image processing techniques such as local feature matching [2] and vegetation density thresholding [33].Ahmadi et al. [2] have presented a visual navigation framework for row crop fields. The system only uses onboard cameras for navigation without performing explicit localization. The switching from one row to another is also implemented in the same framework. They have identified that the agricultural environments are lacking distinguishable landmarks to support localisation and mapping, which causes high visual aliasing. They also state that the constantly growing crops make it difficult to maintain a fixed map of the environment. Their robot could: autonomously navigate through row crop fields without maintaining any global reference maps, and monitor crops in the fields with high coverage by accurately following the crop rows, and the system is robust to fields with various row structures and characteristics. The robotic platform was equipped with two cameras (front and back) to perform the crop row switching. The row-switching method employs feature matching to traverse towards the next crop row. Their work indicates that the vision-based navigation in crop rows could be performed without the need for a global map. The importance of the field of view (FOV) of the cameras used in vision-based navigation systems was indicated by Xue et al. [33] their work on crop row navigation. The researchers used a single monocular camera with a variable FOV setup for better accuracy in navigation at the end of crop rows. Thus the researchers identified the requirement of a wider field of view towards the success of vision-based crop row navigation. The lack of research in fully vision-based crop row switching approaches with vision-based EOR detection indicates a clear gap in the literature within the vision-based infield navigation domain. The four most common headland turning patterns are semi-circular turn, U-turn, light bulb turn (a.k.a $\Omega$ turn) and switch-back turn [10]. The semi-circular turn describes a half circle with a constant radius while the U-turn describes two-quarter circle turns ($90^{\circ}$) centred with a linear traversal stage. These methods are typically used in smaller robots with tighter turning radii relative to the inter-row distance [12, 1, 9]. The $\Omega$ turn and the switch-back turn are mostly used in robots with constrained manoeuvrability such as Ackerman steering robots or tractors [10, 30, 8]. Considering the skid-steering robot configuration of the robotic platform used in this work, the U-turn pattern was chosen to execute the row-switching manoeuvre.
3 Vision-based Navigation in Arable Fields
The vision-based navigation pipeline we propose only uses RGB images for in-row navigation and it uses RGB-D images for row switching. This system does not rely on any assumptions related to the environment it’s deployed, such as type of crop, inter-row space, plant size, sunny or overcast conditions, soil texture and other field variations identified in [28]. An outline of the hardware platform used and the generalization ability of the proposed system to other robotics platforms is presented below. The overall navigation paradigm is also presented in this section.
3.1 Hardware Requirements and Interoperability
The proposed system completely relies on RGBD data from a stereo camera. The camera was mounted onto a mobile robotic platform while its principle axis lies in the $Q2$ quadrant of the vertical plane along the centre of the robot as illustrated in Figure 1. The crop row detection algorithm remains unaffected by the camera oscillations within the $Q2$ quadrant. This navigation system demands a robotic platform which provides wheel odometry. Platforms which provide IMU-corrected wheel odometry provide better performance in crop row switching steps within uneven headland areas.
The proposed system could be deployed on any mobile robotic platform employing an RGBD camera with correct placement as explained above. The crop row detection algorithm could be tuned for optimal operation by adjusting the pitch angle of the camera within $Q2$ followed by the suggested camera placement. A calibration program will lay out virtual guidelines on the camera image frame for this adjustment. A detailed calibration guide on camera placement is presented in our previous work [27]. This system was deployed on two robotic platforms: Clearpath Husky and Hexman Mark-1. However, Hexman Mark-1 robot was used for the experiments outlined in Section 4 due to its longer battery life. Figure 2 shows the robotic setup with a front-mounted Intel RealSense D435i camera and a Reach RS+ RTK GNSS receiver. An NVIDIA Jetson AGX Orin developer kit was used as the onboard computer for the robot. Mark-1 is a skid steer robot with a ground clearance of 128 mm. The dimensions of the robot are $526\times 507\times 244$ mm.
Figure 1: Camera positioning for Triangle Scan Method. The principal axis of the camera(blue) must always reside within Q2.
Figure 2: Hexman Mark-1 robot in the Sugar Beet Field.
3.2 Infield Navigation Scheme
The infield navigation scheme comprises two main components in its navigation behaviour. In-row navigation component [27] is a reactive navigation strategy which uses RGB images from a single front-mounted camera. The captured images are used to detect the angle and linear offset of the central crop row relative to the robot. The angular velocity of the robot is controlled based on these central crop row parameters. The row switching component introduced in Section 3.5 detects the end of the crop row (EOR) and re-entry point to the next row based on RGB images. An aligned depth map of the RGB image was used to identify the distance offset to the next row. This distance offset and EOR position is used in a path planner which executes a U-turn manoeuvre in the form of a 7-step state machine. Figure 3 illustrates the process of the entire infield navigation scheme. The in-row navigation regions are marked with green arrows and the row switching manoeuvre is illustrated in blue.
Figure 3: Vision-based infield navigation scheme. Green: In-row navigation behaviour, Blue: Row switching behaviour.
This infield navigation scheme combines four technical components: a vision-based crop row detection method, a crop row following algorithm based on the detected crop rows, a row switching manoeuvre which perpetuates the row following behaviour culminating an infield navigation scheme and a initial turning direction detection algorithm to initiate the navigation direction. Each of these components are elaborated in the following sections.
3.3 Crop Row Detection
The crop row detection sub-component of this algorithm uses the U-Net [23] based semantic segmentation to generate a crop row mask of the RGB image received from the robot camera. The crop row mask predicted by the U-Net CNN is characterised by a skeleton representation of the crop row rather than attempting to segment the entire region of the crop row within the image as shown in Figure 4. This skeleton representation enables the model to generalize to most of the crops without explicit crop-specific training of the CNN model.
The triangle scan algorithm (TSM) [28] is a post-processing algorithm for the predicted crop row mask which extracts the central crop row line parameters(position and orientation). The TSM algorithm identifies the horizontal position of the vanishing point for all crop rows within the crop row mask by analyzing a narrow strip of pixels at the top of the image. An isosceles triangle ROI is defined with this point being the top position of the triangle and the other two points residing on the bottom edge of the image. The bottom position of the central crop row is scanned within this triangle. These ROIs are illustrated in Figure 5. Angular position ($\Delta\theta$) and displacement ($\Delta L_{x2}$) error extracted by the TSM was used for crop row following.
Figure 4: The proposed crop row following architecture with U-Net CNN for crop row mask detection. The crop mask generated by U-Net CNN is used by a triangle scan method to predict a central crop row ($\Delta\theta$: Crop row angle error corresponding to vertical axis, $\Delta L_{x2}$: Positional error of the central crop row relative to image midpoint). [28]
Figure 5: Regions of interest for scanning the top and bottom points of central crop row. Vanishing point scan ROI: RED, Bottom point scan ROI: Green, H: Height of the image, h: Vanishing point scan ROI height. [28]
3.4 Crop Row Following
The output of the TSM algorithm is used as the input parameters for a robotic controller which steers the robot while it straddles the crop row as shown in Figure 3. A proportional, integral and derivative (PID) controller was used with platform-specific tuning for the robotics platform used in our experiments. This controller stage is considered as a modular entity which can be replaced with any other desired controller. The bipartite crop row error characterised by the TSM is integrated to a single error value $E$ by calculating a weighted sum of the error terms with weights $w_{1}$ and $w_{2}$ as described in Equation 1. This composite error $E$ could be fed into a desired robotic controller to generate an angular velocity signal $\omega$ which would directly control the robot heading relative to the crop row. This control paradigm is best identified as an image-based visual servoing (IBVS) technique [7] where the robot is controlled to minimize an error directly extracted from an image.
${E}=w_{1}\Delta\theta+w_{2}\Delta L_{x2}$ (1)
The crop row following algorithm was executed in a real sugar beet field with varying initial angular errors ranging from -13.53 to 33.14. The tuned PID controller could bring down the error to 80% to 90% of the initial error within 4 meters of traversal into the crop row. The long-distance navigation experiment outlined in Section 4.1 will investigate the accuracy of a row following over long distances while navigating crop rows in a real sugar beet field.
3.5 Crop Row Switching